3. The top four firms in Industry A have market shares of 30, 25, 10, and 5 percent respectively. The top four firms in industry B have market shares of 15, 12, 8 , and 4 percent respectively. Calculate the four-firm concentration ratios for the two industries. Which industry is more concentrated?
Answer #1

Four firms concentration ratio = C = wi

For Industry A

W = maket share

Four firm concentration ratio = C = 30+25+10+5 = 70

For Industry B

Four firms concentration ratio = 15+12+8+4 = 39

Industry A is more concentrated the Industry B